OpenStudy (steve816):

Please help on some trigonometry, it would be greatly appreciated! How to solve this equation?

OpenStudy (steve816):

\[\cos(\theta)+\sin(\theta)=0\]

OpenStudy (steve816):

\[\cos(\theta)+\sin(\theta)=0\]

OpenStudy (love.bug):

rearrange it. Cos theta is what in terms of sine?

OpenStudy (love.bug):

Divide by costheta You get Sin(theta)/Cos(theta) + Cos(theta)/cos(theta) =o [Sinx/cosx = tanx (a law of trigs)] so the same as tan(theta) + 1=0 So tan(theta)=-1

OpenStudy (love.bug):

hope this helps :)